# CosForce: A Force-Based General Pedestrian Model with Anticipation and Reaction Mechanisms

### v3

@@ 3 Numerical Validation
3.1 single-file pedestrians
First, we simulated the single-file motion under periodic boundary conditions, with the variable parameter $\phi$ set to $\pi/3$, as illustrated in Fig.6. To compare the differences in model performance between $\alpha=0$ and $\alpha=0.5$, we set $N=$ $10$, $20$, $30$ and $40$ in the simulation to observe the dynamics of single-file pedestrians under varying densities. The pre-simulation period was set to 500 steps, and simulation data for time steps 501-2300 were analyzed after the pre-simulation (corresponding to 60 s). Since random noise was excluded, the simulation results were deterministic based on the initial conditions. The results for $\alpha=0$ and $\alpha=0.5$ are shown in Fig.6. From the perspective of stability, when $\alpha=0.5$, the model demonstrated superior performance. During the observation period, the traffic flow exhibited synchronization characteristics and no stop-and-go waves were observed at $N=40$. This stability feature aligns with the advantages of the FVD model over the OV model. Corresponding to the angle $\theta$ between the relative speed and the relative distance, we had analyzed the headway-speed relationship, given a sampling frequency of 3 Hz. In the process, two patterns of single-file motion were distinguished: deceleration when $\theta=0$ (space contraction) and acceleration when $\theta=\pi$ (space expansion). When $\alpha=0.5$ (1-D FVD model), a clear hysteresis loop had been observed. In contrast, when $\alpha=0$ (1-D OV model), different motion processes (acceleration, deceleration, synchronized flow) had followed a stable relationship (constant time headway).
@@
The data for empirical validation were obtained from trajectory observations in the single-file experiment, with the details of each experiment provided in Appx.A. Similar to the analysis in Fig.7, the distribution of headway versus speed were examined for different values of $\theta$, given a sampling frequency of 2.5 Hz. The surface fitting formula is expressed by Eq.17, which represents a linear negative correlation function derived from Eq.9. From the fitting results (Fig.8), it can be observed that as the global density increases (situation of $N$=10-45), $\alpha$ approaches zero. This phenomenon indicates the increase in global density resulting in synchronized motion, leading to negligible variations in speed. Consequently, the pedestrian dynamics transition from the FVD model to the OV model.
Figure 7: The relationship between relative distance and speed corresponding to $\theta$: (a) $\alpha=0$, (b) $\alpha=0.5$.
$v_{i}=v_{\text{max}}\left(\tanh\left(\frac{d_{ij}-r_{ij}}{t_{h}}\right)\left(1+\alpha\cos\theta\right)-\alpha\cos\theta\right)$ (17)
Please note that, the surface fitting based on experimental data yields a negative value for $\alpha$, which differs from our setting in the methodology section where $\alpha\in[0,1]$. This discrepancy arises due to the acceleration and deceleration pattern causing an uneven distribution of speed across the range of $\theta$, as observed in Fig.7(b). For this reason, the fitting results could not provide empirical guidance for the calibration of $\alpha$ in the model. However, the observed speed distribution pattern was consistent with the results presented by the model, as shown in Fig.7.
Figure 8: Relationship between relative distance and speed corresponding to $\theta$. The data exhibit a clear trend, with the fitted values of $\alpha$ approaching zero as the density increases.
@@
Fig.9(d) shows the relationship between pedestrian speed and the relative distance to the nearest neighbor. Specifically, based on the constant time headway property, speed exhibits a linear relationship with the nearest neighbor distance under constrained conditions. The distribution trend aligns with empirical findings. A comparison of the data fitting results between experiments and simulations reveals that the experimental time headway (0.804 s) is lower than the simulated result (1.497 s). This discrepancy stems from parameter settings, as the simulation uses a fixed $t_{h}$ of 1.3 s throughout the study.
@@
Fig.9(e) shows the relationship between pedestrian angular velocity and speed, with the constraint exponent of maximum angular velocity corresponding to speed approximately equal to -1.3, slightly lower than our empirical observation (where the constraint exponent is approximately -0.8).
@@ 4 Simulation of "Phase separation" in crowds
Walking and interacting with others are activities that everyone engages in daily. Therefore, preserving intuition is a crucial aspect when investigating crowd dynamics. With this consideration, we have focused on observing basic parameters in statistics, excluding any composite parameters. In crowd dynamics, "phase separation" refers to the phenomenon in which separation occurs between groups with different directions or states, induced by spontaneous order. Common examples include lane formation, stripe formation, and cross-channel formation (Wang, 2024), among others. However, effective metrics for quantitatively estimating self-organization phenomena in crowds are still limited. Based on the mechanisms of self-organization (entropy decrease and steady state), the average normalized speed $\langle v\rangle$, the order parameter $\Phi$, and the average normalized velocity $\langle\mathbf{v}\rangle$ were defined, as given in Eq.19. These metrics serve as quantitative benchmarks for evaluating three distinct aspects of crowd: individual efficiency, system entropy (negatively correlated with the order parameter in crowds), and collective motion at the crowd scale. In the simulation of pedestrian self-organization, since the desired direction is predefined, the order parameters for pedestrians with different directions were calculated independently and averaged to avoid mutual cancellation.
$\langle v\rangle&=\frac{1}{Nv_{max}}\sum\limits_{i\in N}\lVert\mathbf{v}_{i}\rVert&&\text{(average normalized speed)}\\ \Phi&=\frac{1}{N}\left\|\sum_{i=1}^{N}\frac{\mathbf{v}_{i}}{\lVert\mathbf{v}_{i}\rVert}\right\|&&\text{(order parameter)}\\ \langle\mathbf{v}\rangle&=\frac{1}{Nv_{max}}\sum_{i\in N}\mathbf{v}_{i}&&\text{(average normalized velocity)}\end{aligned}\right.$ (19)
@@ 4.3 Cross-channel formation
[…]
Figure 14: Illustration of the cross-channel formation phenomenon. (a) Simulation snapshot. (b) Pedestrian trajectory patterns during the simulation. (c) Formation of sandbars induced by solid-liquid flow interaction. Source: Google Earth, location: 35°13’07"N 93°55’09"E.
@@ 5 Simulation of mass gathering
Self-organization phenomena demonstrate the order and vitality exhibited by a crowd as an organic system. However, this is not the entirety of crowd dynamics. In some cases, crowds may exhibit characteristics entirely opposite where their motion become highly chaotic and locally coherent mechanisms emerge (speed-speed correlation, induced by pushing and shoving), leading to destructive outcomes. As a typical complex system, the simulation of crowd collapse is regarded as a critically important objective of modeling efforts. Specifically, it aims to apply models to simulate crowd accidents and uncover insights that are difficult to obtain through empirical investigations. This represents a critical aspect for understanding the underlying dynamics of crowds and managing them effectively, though it may also be the most challenging aspect.
The continuous dynamism of crowds is one of the fundamental causes of crowd crush. Some characteristics of crowd oscillation have been empirically observed, with contributions from pioneering work (Gu et al., 2024; Ma et al., 2013; Echeverría-Huarte et al., 2022) on the phenomenon of crowd quakes. Before proceeding with the simulation and analysis, it is essential to delineate the limitations of the simulations presented in this paper. As a decentralized system closely intertwined with our daily lives, the dynamics exhibited by crowds are characterized by diversity, stemming from their self-driven and social nature. A most intuitive observation is that the dynamics of crowds vary significantly under different circumstances. This variation is associated with the scale of the crowd, the environment, and the intensity of self-driven behavior induced by social objectives. Among these, the truly dangerous situations arise in dynamical dense crowds, which have prompted lots quantitative analyses based on the properties of velocity fields (Feliciani & Nishinari, 2018; Zanlungo et al., 2023). The crowd system simulated in this paper focuses on the crowd with dynamism property, while static crowds are not considered in this discussion. Our analysis will commence with a specific phenomenon within crowds, namely the "catfish effect," to investigate the impact of dynamic pedestrians on the system of crowd context. Subsequently, we will discuss more generalized crowd dynamics and explore the critical transition processes in dense crowds.
@@
The "catfish effect" in crowd context can be defined as the increase in the overall dynamism of the crowd induced by a minority of dynamic pedestrians, quantitatively marked by the system’s speed gain exceeding the speed increment of dynamic pedestrians. When dynamic pedestrians enter the crowd, adjacent static individuals will gradually adjust their velocity, and the random motion becomes more organized (i.e., polarized). The symmetry breaking resulted in a significant increase in the locally speed. At the microscopic level, this mechanism can be understood as the local polarization of crossing individuals, which has been discussed in lots experiments and models (Bonnemain et al., 2023; Nicolas et al., 2019). This paper focuses on the impact of dynamic pedestrians within the crowd scale and explores the effect phenomena of symmetry breaking and symmetry restoration.
[…]
5.1.1 Normal crowd: interaction force dominated
Based on the parameter settings described above, the ’catfish effect’ was first examined in normal crowds. The crowd size is set as 500, corresponding to a space of 100 m2, with a global density of 5 ped/m2. Such a crowd density resembles that of a typical gathering, where pedestrians can maintain some distance to avoid physical contact. Consequently, the evolution of the crowd is primarily governed by non-contact interaction forces (self-propelled).
Using a simulation without dynamic pedestrians as the baseline, we conducted comparative analysis with unidirectional dynamic pedestrian intrusion (8 dynamic pedestrians), bidirectional dynamic pedestrian intrusion (8 dynamic pedestrians, configured as 4:4), and cross-directional dynamic pedestrian intrusion (8 dynamic pedestrians, configured as 2:2:2:2). A clear illustration is provided in Fig.16(a). Based on the directional configuration of dynamic pedestrians, the simulations are labeled as Normal-0, Normal-1, Normal-2, and Normal-4, respectively. Figs.16(b) and (c) show the distributions of the local density field and velocity field at the moment of 50s in simulation. The local density field reveals the impact of dynamic pedestrians on the spatial variation, where dynamic pedestrians induce heterogeneity in the spatial distribution. Locally, low-density zones form behind the dynamic pedestrians, which is one of the reasons for the polarization phenomena, i.e., the spatial induction of polarization. The velocity field clearly demonstrates these phenomena induced by dynamic pedestrians, leading to an overall speed increase in the crowd. Fig.16(d) presents the cumulative trajectory distribution, with the randomly selected static pedestrians highlighted to observe the movement patterns crowd. In the Normal-1 simulation, the crowd is highly polarized, and the movement of static pedestrians is almost unidirectional. In contrast, in Normal-2 and Normal-4, the polarizing trend caused by dynamic pedestrians is mutually canceled out, and the trajectories show similarities with the configuration of the baseline group.
Based on the statistics of the mean values, the introduction of dynamic pedestrians in simulation Normal-1 indeed induced an increase in the crowd speed compared to the baseline group, as shown in Fig.17. Given the ratio of static to dynamic pedestrians (492:8), the normalized speed gain can be calculated to be approximately 9.2 units, which exceeds the maximum possible speed increment (less than 8 units) that could be attributed to the introduction of 8 dynamic pedestrians. In the simulations of Normal-2 and Normal-4, the speed gains were lower than in Normal-1, thus precluding a quantitative assessment of the "catfish effect". The variation of order parameter indicate that in Normal-1, the introduction of dynamic pedestrians led to a disruption of symmetry within the crowd system, with the polarization trend among the crowd resulting in the "catfish effect". Conversely, in simulations Normal-2 and Normal-4, the introduction of dynamic pedestrians actually enhanced the system’s symmetry (as evidenced by smaller order parameters compared to the baseline group).
Figure 16: Simulation of the "catfish effect" phenomenon in normal crowds. (a) Snapshots of the simulation at 50s: arrows and colors indicate the direction of dynamic pedestrians. (b) The local density distribution at 50 s, with colors representing different types of pedestrians. (c) The velocity field distribution at 50 s, where dots denote the positions of dynamic pedestrians. (d) Trajectories over the 100s simulation period, with bold trajectories highlighting the motion of randomly selected static pedestrians.
Figure 17: Time series variations of the mean normalized speed (a) and the order parameter (b) across different normal crowd configurations.
5.1.2 Dence crowd: contact force dominated
We further investigate the impact of dynamic pedestrians in high-density crowds. The simulated crowd size is set to 1000, corresponding to a simulation space of 100 m2, with a global density of 10 ped/m2. At such a density, pedestrians are tightly packed together, and as a consequence, the interactions within the crowd are primarily governed by contact forces.
Similar with the normal condition, we conducted identical simulations in a dense crowd based on different directional configurations of dynamic pedestrians. The simulations were labeled as Dense-0, Dense-1, Dense-2, and Dense-4, respectively. Fig.18(a) presents snapshots of the four simulation groups at the moment of 50 s. Figs.18(b) and (c) show the corresponding distributions of the local density field and velocity field. In the simulation results, compared to the baseline group (Dense-0), the introduction of dynamic pedestrians in Dense-1 led to a significant increase in crowd velocity, while the speed changes in Dense-2 and Dense-4 were less pronounced. The distribution patterns of the trajectories further support this observation, as illustrated in Fig.18(d).
The results of the quantitative data exhibit a trend consistent with that observed in the Normal crowd, as shown in Fig.19. In the simulation of Dense-1, the speed gain is approximately 16.3 units, which exceeds double the maximum possible speed increment (less than 8 units). However, in the Dense-2 and Dense-4 simulations, the mechanism remains unclear. The data of the order parameter indicate that the polarization trend in Dense-1 is still the highest, while Dense-2 and Dense-4 are very close to each other, both below the baseline group (Dense-0).
@@ 5.1.2 Dence crowd: contact force dominated
Figure 19: Time series variations of the mean normalized speed (a) and the order parameter (b) across different dense crowd configurations.
A horizontal comparison of Fig.17 and Fig.19 reveals that the system’s order parameter is generally higher in dense crowds, indicating a stronger polarization tendency and potentially higher dynamism. This observation is further supported by comparing the average normalized velocity data from the two sets of simulations, see as Fig.20. Compared to normal crowds, the trajectories of dense crowd generally cover a larger range, suggesting more pronounced polarization tendencies and higher velocities.
Figure 20: Time series variations of the average normalized velocity across different configurations in normal and dense crowds. The trajectories in the subplots illustrate the motion pattern of the crowd.
@@ 5.2 Crowd collapse: from crowd quake to crowd avalanche
In section 5.1, in addition to exploring the potential "catfish effect" within the crowd, the higher order parameter in dense crowds have been observed. This perspective illustrate that the increase in crowd density may lead to the process of symmetry breaking in the system and hence may related to crowd collapse.
@@
To simulate crowd collapse and investigate the underlying mechanisms of critical transitions, we established a simulation environment where the global density increases incrementally over time. The settings and parameters, consistent with the simulation in section 5.1, consist of only static pedestrians for generality. In the simulation, the initial crowd density is 5 ped/m2. To create a continuously increasing density scenario for time series observation, we add one pedestrian to the system every 30 time steps. The total simulation duration is 18,000 time steps (corresponding to 600 seconds), during which the maximum density will reaches 11 ped/m2, representing the theoretical upper limit of crowd.
@@
The time series curves of the order parameter and average normalized speed are shown in Fig.21. From the trend of the order parameter, the state transition of the crowd can be identified. The first transition occurs near $\rho=9.8\,\text{ped/m}^{2}$. When the pedestrian density exceeds this threshold, the order parameter begins to rise from a state of steady fluctuations. This phase transition shows a second-order phase transition, where the space between pedestrians is compressed to a delicate equilibrium point. At this point, the crowd is densely packed, and the anticipation and reaction mechanisms completely fail ($\delta=0$). The motion of the crowd shifts from a self-driven process (constrained by the fundamental diagram) to the process of granular dynamics (action equals reaction). The second transition occurs near $\rho=10.4\,\text{ped/m}^{2}$, where the order parameter undergoes a sudden transition, indicating a first-order phase transition. At this moment, the crowd transitions from a highly compressed state to a collapsed state exhibiting chaotic characteristics. Similar transition mechanisms were not observed in the variation of average normalized speed, which transition occurs near $\rho=10.56\,\text{ped/m}^{2}$, approximately 16 s later than the first-order phase transition in the order parameter. The subplot in Fig.21 shows the temporal trend of the average normalized velocity, revealing that the motion of the crowd also undergoes a phase transition similar to that of the order parameter. The trajectory plot shows the crowd’s collective motion in the three phases: low polarization $\rightarrow$ high polarization $\rightarrow$ low polarization with chaos.
@@
Figure 21: Time series variation of the order parameter and average normalized speed. The terms "crowd quakes" and "crowd avalanche" in the figure were adopted from Ma et al. (2013) and Feliciani et al. (2022). For further explanation of the space-speed time delay ($\delta$) in the figure, please refer to Wang et al. (2024).
[…]

### v4

@@ 3 Numerical Validation
3.1 Single-file Dynamics
First, we simulated the single-file motion under periodic boundary conditions, with the variable parameter $\phi$ set to $\pi/3$, as illustrated in Fig.6. To compare the differences in model performance between $\alpha=0$ and $\alpha=0.5$, we set $N=$ $10$, 2020, $30$ and $40$ in the simulation to observe the dynamics of single-file pedestrians under varying densities. The pre-simulation period was set to 500 steps, and simulation data for time steps 501-2300 were analyzed after the pre-simulation (corresponding to 60 s). Since random noise was excluded, the simulation results were deterministic based on the initial conditions. The results for $\alpha=0$ and $\alpha=0.5$ are shown in Fig.6. From the perspective of stability, when $\alpha=0.5$, the model demonstrated superior performance. During the observation period, the traffic flow exhibited synchronization characteristics and no stop-and-go waves were observed at $N=40$. This stability property aligns with the advantages of the FVD model over the OV model.
@@ 3.2 Unidirectional Flow
Fig.7(d) shows the relationship between pedestrian speed and the relative distance to the nearest neighbor. Specifically, based on the constant time headway property, speed exhibits a linear relationship with the nearest neighbor distance under constrained conditions. The variation trend aligns with empirical findings. A comparison of the data fitting results between experiments and simulations reveals that the experimental time headway (0.804 s) is lower than the simulated result (1.497 s), with the fitting function given by Eq. 17. This discrepancy stems from parameter settings, as the simulation uses a fixed $t_{h}$ of 1.3 s throughout the study.
@@
Fig.7(e) shows the relationship between pedestrian angular velocity and speed, with the constraint exponent of maximum angular velocity corresponding to speed approximately equal to -1.3, slightly lower than our empirical observation (where the constraint exponent is approximately -0.8 (Wang et al., 2025)).
@@ 4 Simulation of "Phase separation" in crowds
Walking and interacting with others are activities that everyone engages in daily. Therefore, preserving intuition is a crucial aspect when investigating crowd dynamics. With this consideration, we have focused on observing fundamental parameters in statistics, excluding any composite parameters. In crowd dynamics, "phase separation" refers to the phenomenon in which separation occurs between groups with different directions or states, induced by spontaneous order. Common examples include lane formation, stripe formation, and cross-channel formation (Wang, 2024), among others. However, effective metrics for quantitatively estimating self-organization phenomena in crowds are still limited. Based on the mechanisms of self-organization (entropy decrease and steady state), the average normalized speed $\langle v\rangle$, the order parameter $\Phi$, and the average normalized velocity $\langle\mathbf{v}\rangle$ were defined, as given in Eq.18. These metrics serve as quantitative benchmarks for evaluating three distinct aspects of crowd: individual efficiency, system entropy (negatively correlated with the order parameter in crowds), and collective motion at the crowd scale. In the simulation of pedestrian self-organization, since the desired direction is predefined, the order parameters for pedestrians with different directions were calculated independently and averaged to avoid mutual cancellation.
$\left\{\begin{aligned} \langle v\rangle&=\frac{1}{Nv_{max}}\sum\limits_{i\in N}\lVert\mathbf{v}_{i}\rVert&&\text{(average normalized speed)}\\ \Phi&=\frac{1}{N}\left\|\sum_{i=1}^{N}\frac{\mathbf{v}_{i}}{\lVert\mathbf{v}_{i}\rVert}\right\|&&\text{(order parameter)}\\ \langle\mathbf{v}\rangle&=\frac{1}{Nv_{max}}\sum_{i\in N}\mathbf{v}_{i}&&\text{(average normalized velocity)}\end{aligned}\right$. (18)
@@ 4.3 Cross-channel formation
The final scenario in the simulation of phase separation is referred to as "cross-channel formation", which describes the phenomenon where pedestrians, when crossing through static crowds, form a series of stable cross-channels. A field video of this phenomenon can be observed in a train station (Wang, 2024). In our simulations, the total number of pedestrians was set as 250, with a proportionally varied pedestrian state configuration (red refer dynamic pedestrian and blue refer static pedestrian). The variable parameters were set as $v_{\rm{max}}=1.4\;\text{m/s}$ (dynamic) or $0\;\text{m/s}$ (static), $\phi=\pi/2$ (dynamic) or $\pi$ (static), and $\alpha=0.5$.
Figure 12: Illustration of the cross-channel formation phenomenon. (a) Simulation snapshot. (b) Pedestrian trajectory patterns during the simulation. (c) Formation of sandbars induced by solid-liquid flow interaction. Source: Google Earth, location: 3513’07"N 9355’09"E.
@@ 5 Simulation of mass gathering
Self-organization phenomena demonstrate the order and vitality exhibited by a crowd as an organic system. However, this is not the entirety of crowd dynamics. In some cases, crowds may exhibit characteristics entirely opposite where their motion become highly chaotic and locally coherent mechanisms emerge (velocity correlation, induced by pushing and shoving), leading to destructive outcomes. As a open challenge, the simulation of crowd collapse is regarded as a critically important objective of modeling efforts. Specifically, it aims to apply models to simulate crowd accidents and uncover insights that are difficult to obtain through empirical investigations. This represents a critical aspect for understanding the underlying dynamics of crowds and managing them effectively, though it may also be the most challenging aspect.
The continuous dynamism of crowds is one of the fundamental causes of crowd crush. Some characteristics of crowd oscillation have been empirically observed, with contributions from pioneering works (Gu et al., 2024; Ma et al., 2013; Echeverr´ıa-Huarte et al., 2022) on the phenomenon of crowd quakes. Before proceeding with the simulation and analysis, it is essential to delineate the limitations of the simulations presented in this paper. As a decentralized system closely intertwined with our daily lives, the dynamics exhibited by crowds are characterized by diversity, stemming from their self-propelled property and social nature. A most intuitive observation is that the dynamics of crowds vary significantly under different circumstances. This variation is associated with the scale of the crowd, the environment, and the intensity of self-propelled behavior induced by social objectives. Among these, the truly dangerous situations arise in dynamical dense crowds, which have prompted lots quantitative analyses based on the properties of velocity fields (Feliciani & Nishinari, 2018; Zanlungo et al., 2023). The crowd system simulated in this paper focuses on the crowd with dynamism property, while static crowds are not considered in this discussion. Moreover, the two-dimensional simulation in this study cannot capture mechanisms such as pedestrian falls, which may lead to distortions in simulating crowd collapse. Therefore, awareness of the model’s limitations is essential. Our analysis will commence with a specific phenomenon within crowds, namely the "catfish effect," to investigate the impact of dynamic pedestrians on the system of crowd context. Subsequently, we will discuss more generalized crowd dynamics and explore the critical transition processes in dense crowds.
@@
The "catfish effect" in crowd context can be defined as the increase in the overall dynamism of the crowd induced by a minority of dynamic pedestrians, quantitatively marked by the system’s speed gain exceeding the speed increment of dynamic pedestrians. When dynamic pedestrians enter the crowd, adjacent static individuals will gradually adjust their velocity, and the random motion becomes more organized (i.e., polarized). The symmetry breaking resulted in a significant increase in the locally speed. At the microscopic level, this mechanism can be understood as the local polarization of crossing individuals, which has been discussed in lots experiments and models (Bonnemain et al., 2023; Nicolas et al., 2019). This paper focuses on the impact of dynamic pedestrians within the crowd scale and explores the effect of symmetry variation.
@@
To simulate these mechanisms, variable parameters were set as ${v}_{max}=1.4\;\text{m/s}$ (dynamic) or $0.6\;\text{m/s}\;[0,0]$ (static), $\phi=\pi/2$ (dynamic) or $\pi$ (static), and $\alpha=0.5$. The zero vector $[0,0]$ represents the direction of the maximum velocity. According to Eqs.11 and 12, under these parameter settings, the self-driven force will try to slow down static pedestrians to a speed of 0, while the repulsive force will be induced by nearest neighbor, scale corresponds to a maximum speed of 0.6 m/s. This setup leads to static pedestrians being highly sensitive to spatial variations and leads to a structured distribution, which is characteristic of dynamic crowds. Overall, the crowd mainly consists of static pedestrians, providing the background environment of the system. Our aim is to study changes in the state caused by dynamic pedestrians and explore the potential phenomenon of the "catfish effect", which originates from our earlier conjecture (Wang et al., 2023). The simulation space is set as an 10 m $\times$ 10 m with fully periodic boundary conditions. The simulated crowd size is set to 1000, corresponding to a simulation space of 100 m2, with a global density of 10 ped/m2. At such a density, pedestrians are tightly packed together, and as a consequence, the interactions within the crowd are primarily governed by contact forces. In the initial state, all pedestrians are randomly distributed, with an initial speed of 0 m/s. The simulation time for a single scenario is 3000 time steps (corresponding to 100 s).
A simulation scenario without dynamic pedestrians served as the control group for comparative analysis against three intrusion types: unidirectional intrusion (8 dynamic pedestrians), bidirectional intrusion (8 dynamic pedestrians, configured as 4:4), and cross-directional intrusion (8 dynamic pedestrians, configured as 2:2:2:2). Based on the directional configuration, the simulations are labeled as Dense-0, Dense-1, Dense-2, and Dense-4, respectively. Fig.14(a) presents snapshots of the four simulation groups at the moment of 50 s. Figs.14(b) and (c) show the corresponding distributions of the local density field and velocity field. In the simulation results, compared to the control group (Dense-0), the introduction of dynamic pedestrians in Dense-1 led to a significant increase in speed, while the speed changes in Dense-2 and Dense-4 were less pronounced. The distribution patterns of the trajectories further support this observation, as illustrated in Fig.14(d).
Fig. 15 shows the time series variations of the mean normalized speed and the order parameter across different crowd configurations. In the Dense-1 simulation, the speed gain reaches approximately 16.3 units, more than twice the maximum possible increment (less than 8 units). However, the underlying mechanism remains unclear in the Dense-2 and Dense-4 simulations. The order parameter data suggest that the polarization tendency in Dense-1 remains the strongest, while Dense-2 and Dense-4 show very similar trends, both falling below the control group (Dense-0). Fig. 16 presents the variation trends of the average normalized velocity under different crowd configurations. It can be observed that in the unbalanced dynamic pedestrian configuration (Dense-1), the crowd exhibits a stronger polarization tendency. In contrast, this difference is not evident in the balanced dynamic configurations. This effect may be attributed to global statistical observations masking the polarization tendency in relative orientations.
Figure 15: Time series variations of the mean normalized speed (a) and the order parameter (b) across different crowd configurations.
Figure 16: Time series variations of the average normalized velocity across different crowd configurations. The trajectories in the subplots illustrate the motion pattern in the crowd scale.
@@ 5.2 Crowd collapse: from crowd quake to crowd avalanche
In Section 5.1, beyond analyzing the potential "catfish effect" within the crowd, we identified variations in the order parameter, indicating that changes in collective symmetry may be related to the onset of crowd collapse. To simulate crowd collapse and investigate the underlying mechanism, we established a simulation scenario where the global density increases incrementally over time. The settings and parameters, consistent with the simulation in section 5.1, consist of only static pedestrians for generality. In the simulation, the initial crowd density is 5 ped/m2. To create a continuously increasing density scenario for time series observation, we add one pedestrian to the system every 30 time steps. The total simulation duration is 18,000 time steps (corresponding to 600 seconds), during which the maximum density will reaches 11 ped/m2, representing the theoretical upper limit of crowd density.
@@
The time series curves of the order parameter and average normalized speed are shown in Fig.17(a). From the trend of the order parameter, the state transition of the crowd can be identified. The first transition occurs near $\rho=9.8\,\text{ped/m}^{2}$. When the pedestrian density exceeds this threshold, the order parameter begins to rise from a state of steady fluctuations. This transition shows a second-order transition, where the space between pedestrians is compressed to a delicate equilibrium point. The motion of the crowd shifts from a self-propelled process (constrained by the fundamental diagram) to the process of granular dynamics (action equals reaction). The second transition occurs near $\rho=10.4\,\text{ped/m}^{2}$, where the order parameter undergoes a sudden transition, indicating a first-order transition. At this moment, the crowd transitions from a highly compressed state to a collapsed state exhibiting chaotic characteristics. Transition of average normalized speed shows a delay, which occurs near $\rho=10.56\,\text{ped/m}^{2}$, approximately 16 s later than the transition in the order parameter. The subplot in Fig.17(a) shows the temporal trend of the average normalized velocity, revealing that the motion of the crowd also undergoes similar transitions. The trajectory plot shows the crowd’s collective motion in the three states: low polarization $\rightarrow$ high polarization $\rightarrow$ low polarization with high dynamism.
The critical transition process can be intuitively explained through the analogy of a tipping balance scale, as shown in Fig.17(b). At low densities, the crowd system exhibits adaptive behavior. Its state remains largely unaffected by environmental influences. As both density and dynamism increase, the crowd tends to a more compact and structured distribution. This is accompanied by a decline in robustness and tendency of polarization. Under such conditions, the crowd exists in a delicate equilibrium, analogous to a balanced scale, where even slight perturbations can induce a transition. The transition from equilibrium to deflection is similar to the propagation of local imbalance among individuals in the crowd. This imbalance gradually spreads and evolves into a domino effect. Such a process accounts for the delayed onset of speed transition observed in the simulation.
@@
Figure 17: Critical transition process of crowd system. (a) time series variation of the order parameter and average normalized speed, (b) balance scale illustration of the critical transition in crowd dynamics. The terms "crowd quakes" and "crowd avalanche" were adopted from Ma et al. (2013) and Feliciani et al. (2022), $\varepsilon$ denotes the perturbation.
Critical moments observed during the simulation correspond to the times marked by vertical lines in Fig.17(a), with densities of 9, 9.8, 10.4, 10.56, and 10.8 ped/m2, respectively. We present the simulation scenarios along with the corresponding density fields, velocity fields, and the field of divergence and vorticity, as shown in Fig.18. At the moment corresponding to $\rho=9$ ped/m2, the crowd remains in a stable state. Pedestrians exhibit a relatively uniform spatial distribution. When the density increases to 9.8 ped/m2, spatial contraction among individuals and an increase in pedestrian speed can be observed. This corresponds to the moment of the second-order transition in Fig. 17. As the density further rises to 10.4 ped/m2, the system reaches the first-order transition point. At this stage, pedestrian compression approaches a critical limit, beyond which further compression becomes nearly impossible. Mutual physical forces and motion interactions render individuals almost unable to maintain balance. This imbalance initially occurs at the individual level and indicates a disruption of force equilibrium. In real-world scenarios, such a condition may lead to pedestrian falls. This imbalance breaks the stability of the crowd system and marks the tipping point when the crowd begins to shift. The moment $\rho=10.56$ ped/m2 corresponds to the process of crowd collapse, as shown in Fig.18(d), where a domino-like propagation mechanism is clearly visible. The immense internal energy generated by crowd compression erupts upon reaching the compression limit, resulting in the occurrence of crowd collapse. Fig.18(e) demonstrates the crowd collapse phenomenon observed in the simulation. The periodic boundary conditions and the exclusive consideration of force transmission overlook potential falling processes during crowd collapse. This leads to a notable discrepancy from real scenarios.